### v5

# Non-Smooth Solutions of the Navier-Stokes Equation

Abstract
Non-smooth Leray-Hopf solutions of the Navier-Stokes equation are constructed. The construction occurs in a finite periodic cube $\mathbb{T}^{3}$. The entropy production maximizing solutions are selected.
Part I of this paper defines the entropy principle and using it, finds improved regularity for the Navier-Stokes solutions.
Part II concerns initial data and its achievability as a limit of the small time data.
Part III establishes analyticity properties of the Part II solution; Part IV demonstrates blowup in finite time.
I Introduction
I.1 Main Results
The main result of this paper is the construction of non-smooth entropy production maximizing solutions of the Navier-Stokes equation of the Leray-Hopf (LH) class of weak solutions.
The proof is divided into four main steps, given in Parts I, II, III, IV of this paper.
Part I formulates the fixed time entropy principle, and assuming this principle, establishes improved regularity of the Navier-Stokes solution.
Part II demonstrates the achievability of the initial data from limits of the small time data.
Part III establishes analyticity of the weak solutions of the Navier-Stokes equation.
Part IV proves blowup in finite time for initial conditions constructed with the help of Parts II and III.
In this paper, only single incompressible fluids (not mixtures) are considered. The fluids are real (not complex) valued. They are defined in a periodic spatial domain $\mathbb{T}^{3}$.
I.2 Prior Results
A review of Navier-Stokes solutions can be found in BedVic22 . The criteria of BeaKatMaj84 are necessary for blowup. The numerical study Hou23 with additional numerical and theoretical references cited provide strong evidence for the existence of non-smooth solutions of the Navier-Stokes equation.
Weak solutions for the Euler equation have been constructed, with weakly defined limits and function spaces DiPMaj87 ; BreDeLSzk11 ; BreFei18 ; Neu93 ; GanWes09 ; BucVic19 ; KukVicWan20 and additional references cited in these sources. See Ber23 and references cited there for a discussion of the Navier-Stokes and Euler equations and their relation to the Onsager conjecture.
The Foias Foi74 ; FoiManRos04 ; RobRodSad16 description of Navier-Stokes turbulence as an LH weak solution provides the constructive step needed in our analysis. Other ideas, including analyticity, from this body of work are also used.
The existence of entropy production maximizing solutions of the Navier-Stokes equation is established in GliLazChe20a ; CheGliSai24b
The theory of vector and tensor spherical harmonics VarMosKhe88 is used.
The Lagrangian for space time smooth fluids, derived from Jon15 ; FukFuj10 ; FukFuj12 ; HolMarRat98 ; Tay11 ; SelWhi68 , has influenced our thinking. Scaling analysis, based on the renormalization group has also influenced our thinking. Of the many papers on this subject, we mention BogShi59 ; MonYag75 ; Wil83
Additional references are added as needed throughout the text. Additional references can be found within the sources cited above. This paper is not a scholarly review.
II Preliminaries
II.1 Function Spaces
Configuration Space
[…]
The bare vacuum energy of the fluid, $\langle u^{2}\rangle$, is a central concern of this paper.
Powers of the Stokes Operator
The Stokes operator is denoted $A$. For space periodic boundary conditions,
$A=-\nu\Delta\ ,$ (1)
where $\Delta$ is the Laplace operator, FoiManRos04 , Ch. II.6, Eq. (6.16).
The fractional power of the Stokes operator is $(-\nu\Delta)^{s}$. Using these powers, we define fractional Hilbert spaces $\mathcal{H}_{s}$ with the inner product $(u,v)_{\mathcal{H}_{s}}=(u,(-\nu\Delta)^{s}v)$. Incompressibility is imposed as part of the definition of $\mathcal{H}_{s}$.
The Spaces $\mathcal{V}$ and $\mathcal{V}^{\prime}$
The space $\mathcal{H}_{1}$ is denoted $\mathcal{V}.$ Its dual is $\mathcal{V}^{\prime}=\mathcal{H}_{-1}$.
II.2 Euler and Navier-Stokes Equations
The Euler and Navier-Stokes equations are defined physically as the law of conservation of momentum. The weak solutions of these equations are interpreted through space time integration, meaning that they are valid as an expectation value with a cutoff function $\phi$, as in (11).
The existence of nonunique and nonphysical solutions of these equations have been known for the 1D Euler equation since the days of Riemann, for the 2D Euler equation Sch93 and in 3D even for the Navier-Stokes equation; see BucVic19 ; BucVic20 and references cited there.
Evaluation of a divergent integral in a Sobolev space of negative index does not change the theory, but modifies the function spaces in which it is described. In view of the infinities of turbulence, these negative index Sobolev spaces are essential and can never be removed. In contrast, a cutoff, if used, is temporary and removed in the sense of limits.
The Leray Projector
Incompressibility, $\sum_{i=1}^{3}\partial_{i}u_{i}=0$, is imposed on all vector fields.
Define the Leray projector $\mathbb{P}$ as the projection onto the divergence free subspace of velocity fields, acting on an $L_{2}$ space.
$\mathbb{P}_{q}$ denotes the Leray projector acting on an $L_{q}$ space, with an adjoint $\mathbb{P}_{p}$ acting on the dual $L_{p}$ space. See Eva10 .
The convective derivative is defined as
$\frac{D}{Dt}=\frac{\partial}{\partial t}+u\cdot\nabla\ .$ (2)
The Euler equation, expressed using Lagrangian derivatives, is
$\mathbb{P}\frac{Du_{i}}{Dt}=f_{i}\ ,$ (3)
with $f$ the stirring force, assumed divergence-free.
The Stirring Force $\mathbf{f}$
Restrictions on $f$ are based on physics, as unrestricted stirring is consistent mathematically with the Euler and Navier-Stokes equations. The choice $f=0$ is made within this paper.
Stress Rates
$S_{ij}$ is the rate of stress tensor, defined as
$S_{ij}=\frac{\partial u_{i}}{\partial x_{j}}\ .$ (4)
Vorticity $\omega$ is defined as
$\omega_{1}=\nabla\times u$ (5)
and the vorticity related rate of stress is
×$\nabla\times S_{ij}=\frac{\partial\omega_{i}}{\partial x_{j}}\ .$ (6)
The stress-strain tensor
$\sigma_{ij}=\frac{1}{2}\Big{(}S_{ij}+S_{ji}\Big{)}$ (7)
specifies the fluids modeled, an incompressible isotropic Newtonian fluid.
The Navier-Stokes equation written using Lagrangian derivatives is
$\mathbb{P}\frac{Du_{i}}{Dt}=\nu\mathbb{P}(\partial_{j}\cdot S_{ij})+f_{i}\ .$ (8)
The Eulerian Navier-Stokes equation,
$\bigg{(}\frac{\partial u_{i}}{\partial t}+\mathbb{P}\Big{(}u_{j}\frac{\partial u_{i}}{\partial x_{j}}\Big{)}\bigg{)}=\nu\mathbb{P}(\nabla_{j}\cdot\sigma_{ij}(x,t))+f_{i}\ ,$ (9)
is formally equivalent to (8) with $\nu=0$ to define the Euler equation.
II.3 Cutoffs
Cutoffs are defined as a product of a spatial and a temporal cutoff. The spatial cutoff is the projection of the velocity field $u$ onto a finite dimensional space of smooth basis elements of $\mathcal{H}$. The temporal cutoff is defined by convolution by a smooth test function, which is time retarded in the sense of being supported in $t<0$.
A cutoff is defined as a modification to the theory. The cutoff theory restricts the spatial integration to a finite dimensional subspace of the configuration space $\mathcal{H}$.
The complete criteria to define a weak solution of the Navier-Stokes equation are formulated in terms of cutoffs. Cutoffs are defined as functions $\phi$ belonging to
$\begin{split}&\widetilde{\mathcal{D}}_{\sigma}(\mathbb{T}^{3})=\Big{\{}\phi\;:\;\phi=\sum_{k=1}^{N}\alpha_{k}(t)a_{k}(x),\\ &\alpha_{k}\in C_{c}^{1}([0,\infty)),\;a_{k}\in\mathcal{N},\;N\in\mathbb{N}\Big{\}}\ .\\ \end{split}$ (10)
Here $\mathcal{C}^{1}_{c}$ denotes the space of continuously differentiable functions with compact support and $\mathcal{N}$ is the set of eigenfunctions of the Stokes operator, as defined in RobRodSad16 , Theorem 2.24. (Caution: $\alpha_{k}\neq a_{k}$ in (10).)
Following RobRodSad16 , Def. 3.3, a weak solution of the Navier-Stokes equation is defined as
  1. 1.
    $u_{0}\in\mathcal{H}$
  2. 2.
    $u\in L_{\infty}(0,T;\mathcal{H})\cap L_{2}(0,T;\mathcal{H}_{1})$
  3. 3.
    $u$ satisfies weak conservation of energy with $\phi\in\widetilde{{D_{\sigma}}}(\mathbb{T}^{3})$ as test functions, according to RobRodSad16 , Eq. (3.1).
    $\begin{split}&\int_{0}^{s}-(u,\partial_{t}\phi)+\int_{0}^{s}\nu(\nabla u,\nabla\phi)+\int_{0}^{s}((u\cdot\nabla u),\phi)\\ &=(u_{0},\phi(0))-(u(s),\phi(s))\end{split}$ (11)
    for a.e. $s>0$.
Consequences of the definition, items 1–3, of a weak solution are given in RobRodSad16 .
By RobRodSad16 Exercise 3.4, $u$ solves the Navier-Stokes equation weakly, relative to test functions in $\mathcal{H}_{1}$. Thus $u$ solves the Navier-Stokes equation in $\mathcal{V}^{\prime}$.
By RobRodSad16 , Lemma 3.4 and 3.7,
$\begin{split}&\partial_{t}u\in L^{4/3}(0,T);\mathcal{V}^{\prime})\\ &(u\cdot\nabla u)\in L^{4/3}(0,T;L^{6/5})\ .\\ \end{split}$ (12)
II.4 Second Order Phase Transitions of Decaying Turbulence
Decay processes, referring to the entire turbulent process and not to individual states within it, are subsumed in the generic notion of a second order phase transition, whose ultimate rest state is the hydrostatic state of a fluid at rest.
The second order phase transition occurs in four main steps, as indicated here in a formal analysis and with the first three steps studied numerically in HsuKauGli22 , based on the analysis of DNS simulation data.
The first step is the decay of enstrophy, from a highly stirred (high vorticity) state to a state of near zero enstrophy. The second step is decay in the energy mode, driven by the cascade of energy states to smaller wavelength, where they ultimately vanish, and ending in the vacuum state of fully developed turbulence. The third step is driven by viscous dissipation, proceeding in a fourth step by radiation of thermal energy to a regime in thermal equilibrium with an ambient temperature and only thermally residual molecular motion.
The second order phase transitions for the Euler equation have the first two and fourth steps, while lacking the viscosity related third step.
II.5 Harmonic Analysis
The rotational and translation groups combine to form a semidirect product. The rotational and translation degrees of freedom of a random field, all but its absolute magnitude, are defined by its invariance properties under this semidirect product.
Fourier Analysis
Harmonic analysis of the translation invariance subgroup of the semidirect product is defined by Fourier analysis, which expands the random field into one dimensional Fourier components.
Any $u\in\mathcal{H}$, the $L_{2}$ space of turbulent fields $u$, has the Fourier expansion
$u=\sum_{k\in\mathbb{Z}^{3}}\hat{u}_{k}e^{2}\pi ix/L\ ,$ (13)
with $L$ the length of one side of the periodic domain, subject to the incompressibility condition
$k\cdot\hat{u}_{k}=0$ (14)
for all $k\in\mathbb{Z}^{3}$. Then
$L^{3}\sum_{{k\in\mathbb{Z}^{3}}}|\hat{u}_{k}|^{2}=|u|^{2}<\infty\ .$ (15)
Vector Spherical Harmonics
Spherical harmonics are described by a sum of representations of the rotation group, $O(3)$. For a scalar random field, the expansion is defined by an expansion in the basis elements $Y_{\ell m}$, with $\ell\in\mathbb{N}$ and $-\ell\leq m\leq\ell$.
Vector spherical harmonics describe the rotational degrees of freedom of a random vector field, such as a fluid velocity field. A possible spin variable to identify the particle transformation law is set to 1 for fluids. In this case, the vector spherical harmonics are described by three representations of $O(3)$. The full theory of tensor spherical harmonics is given in VarMosKhe88 .
II.5.1 Solid Spherical Harmonics
The customary spherical harmonics, with the basis $Y_{\ell,m}$, are augmented with a magnitude $r$. The complete expansion has the basis labeled $r,\ell,m$. On restriction to the unit sphere, $r$ takes on two values, $r=\pm 1$. The plus sign selects the solid spherical harmonics solution that is regular at the origin. The minus sign selects the solid spherical harmonics solution that is regular at infinity. Detailed accounts of spherical and solid mechanics for vector and tensor fields is given in VarMosKhe88 .
Solid mechanics harmonic analysis refers to functions that depend on $r$ as well as on the asmuthal and polar angles $\phi$ and $\theta$ which parameterize the unit sphere in $R^{3}$. MorFes53 ; Noe1918 ; WanRonBur08 ; TouSto77 . The solid mechanics solution $\Psi$, defined in terms of the parameters $(r,\theta,\phi)$ has the radial and angular separation:
$\Psi(r,\theta,\phi)=R(r)Y_{\ell,m}(\theta,\phi)\ .$ (16)
The radial part $R(r)$ has two solution classes: the spherical Bessel and Neumann functions:
$R(r)=j_{l}(kr)+y_{l}(kr)\ .$ (17)
The Neumann functions are excluded by a non-singularity condition at the origin. A singularity occurs only in the Neumann term. The general solutions are:
$\Psi_{klm}^{smooth}(r,\theta,\phi)$ $=$ $\sqrt{\frac{2}{\pi}}kj_{l}(kr)Y_{l,m}(\theta,\phi)$ (18)
$\hat{\Psi}_{klm}(r,\theta,\phi)$ $=$ $\sqrt{\frac{2}{\pi}}ky_{l}(kr)Y_{l,m}(\theta,\phi)$ (19)
Any function $f(r,\theta,\phi)$ has the Fourier expansion WanRonBur08 :
[…]
The expansion of the total energy dissipation rate $\epsilon$ has a special form. The derivation HigKop68 shows that the index $m$ does not occur in the integral,
[…]
The coefficients $C,F,S$ define the $\epsilon$ analysis. Their asymptotics relate to the regularity of $\epsilon$ in the standard way as in Fourier analysis.
Solid mechanics analysis allows a detailed evaluation of the volume integrals that determine the energy levels within a given $l$ index subspace. The combined $r,\ell,m$ analysis defines a complete expansion based on the $O(3)$ orthogonal group.
The $r$ solid harmonics expansion of a polynomial function of $u$ is independent of the spatial coordinate $x$. It is a complete description of the $t$ dependence. The $\ell,m$ expansion is independent of the time coordinate $t$ and it is a complete description of the $x$ dependence. Thus $r$ defines dynamics and $\ell,m$ define kinematics.
Both the $r$ and the $\ell,m$ variables have the dimension of energy. Their entropies are related via a Legendre transform (proven for cutoff theories) to the energy dissipation in the $r$ and $\ell$ variables.
Upon restriction to the unit sphere, the $r$ expansion takes on the values ±$\pm 1$. With this restriction $r$ is dimensionless.
Spherical harmonics extend to the vector fields labeled by energy and $\pm$ enstrophy. This extension applies to solid harmonics, with additional labels for $\Psi$ and $\hat{\Psi}$ to describe the solid mechanics solution.
Fixed Time and Fixed Space Harmonics
The functions $Y_{\ell,m}$ provide a complete orthonormal basis for the spatial (fixed time) $L_{2}$ space, with the restriction $r=1$. The result is a function of the three spatial variables.
General functions are then a product of the fixed space functions and a remainder.
The remainder is the $r$ expansion, defined at a fixed spatial location. It is a function of time. The remainder is described by $r$ as a configuration space variable and by a conjugate momentum $\pi$. $\pi$, defined at a fixed space location, is a function of $x_{1},x_{2},x_{3}$ equivalence classes. $\pi$ thus has a single (generic) spatial variable, which does not refer to any individual $x_{1},x_{2},x_{3}$.
II.5.2 Quotient Spaces
We define a quotient space by identifying all fixed time harmonic expansion terms. The quotient space is isomorphic to the fixed space harmonic expansion.
The fixed time spherical harmonics are defined by restriction of all variables to be spatial only. The result is a function of three spatial variables. From the point of view of the fixed time expansion, $r$ is a scalar. $r$ is eliminated on restriction to the sphere with $r=\pm 1$ and selection of the solution regular at the origin, with $r=1$. At fixed time, all variables are in configuration space, which is defined as functions of 3 spatial variables.
The fixed space spherical harmonics are defined with one temporal variable, a configuration space variable.
In the $r$ expansion, $r$ is the one dimensional configuration space variable. Associated with this variable is its conjugate momenta $\pi$, which is a stochastic function of a one dimensional variable.
Since $r$ and $\pi$ are defined at fixed space variables, the spatially dependent $\pi$ is defined over permutation equivalence classes of spatial variables. $\pi$ is thus a function of time and a generic (1D) spatial variable. The spatial variable has no reference to specific $x_{i},x_{2},x_{3}$ spatial coordinates. $r$ and $\pi$ define loop variables. The loop variables provide a complete basis for the $r,\pi$ expansion, defined at a fixed spatial location.
Young’s Inequality
Young’s inequality FoiManRos04 , eq. (4.17), pg. 20. states that
$ab\leq\frac{a^{p}}{p}+\frac{b^{p^{\prime}}}{p^{\prime}}$ (22)
for all $a,b>0$ and for $p,p^{\prime}$ Holder conjugate indices.
Young’s inequality can be used to bound cross terms, so that the details of the quotient spaces can by bypassed with separate estimates of the fixed space and fixed time cases.
III Part I: The Entropy Principle
The entropy principle as used in this paper states that the solution $u(x,t)$ of the Navier-Stokes equation, maximizes the production of entropy relative to other weak solutions of the Navier-Stokes equation. This principle is applied within the space of solid spherical harmonic expansion coefficients. It is convenient to maximize individually the entropy in both the $r,\pi$ and the $\ell,m$ parameters.
The entropy maximizing state is computed as the equally weighted sum for the $r,\pi$ and $\ell,m$ values.
See GliLazChe20a ; CheGliSai24b for a related discussion. Uniqueness of the entropy maximizing state will be addressed in a future publication with G.-Q. Chen and H. Said.
III.1 Dissipation Operators
We will be concerned with various forms of turbulent and viscous dissipation, as a function of $u$. The notations are given by (23).
$\begin{split}&\nu_{t,Temp}=-\langle u_{i},\partial_{t}u_{k}\rangle\\ &\nu_{t,NL}=\sum_{j}\langle u_{j}\partial_{j}u_{i}\rangle\\ &\nu_{t}=\nu_{t,NL}+\nu_{t,Temp}\\ &\nu_{Tot}=\nu_{t,Temp}+\nu_{t,NL,i},u_{k}+\nu\|\nabla u\|^{2}_{L_{s}}\ .\\ \end{split}$ (23)
Higher moments of the sums in (23) are bounded by the sum of the higher moments of the individual terms. To justify this statement, the cross terms are bounded using Young’s inequality.
Young’s inequality is used pairwise, and first serves to bound viscous vs. turbulent cross terms. The same analysis is repeated for the cross terms between the two turbulent dissipation terms.
Positivity of the $\nu_{t}$ Variables
  • Proposition III.1.1 Entropy maximizing dissipation operators $\nu_{t,Temp}$ and $\nu_{t,NL}$ are non-negative in their energy dissipation.
Proof. We treat entropy maximizing and energy dissipation maximizing as identical. The maximum rate of energy dissipation chooses a positive sign for the dissipation. $\square$
III.2 Conservation of Energy
A weak form of the (exact) conservation of energy is contained in Lee24v6 , with an elegant and simple proof. We note the relation of this proof to earlier studies of Lad63 ; MikSoh88 ; GigSoh91 . The proof assumes a sum over repeated indices (SRI). We adapt it to the case NSRI (which is defined as no SRI). It will play a key role in the paper.
The result of Lee24 is formulated as a Lemma 1:
  • Lemma III.1 Let $u$ be a weak solution of the Navier-Stokes equation. Then
    $\langle\partial_{t}u,u\rangle=-\|\nabla u\|^{2}$ (24)
    for a.e. $t$, where the brackets denote a dual pairing between $\mathcal{V}^{\prime}$ and $\mathcal{V}$.
Proof. We have already observed that the Navier-Stokes equation is satisfied for a.e. $t$ while $u\in\mathcal{V}$. We take the inner product with $u$ and vacuum expectation values to obtain
$\langle\partial_{t}u,u\rangle=\nu\langle\Delta u,u\rangle-\langle(u\cdot\nabla)u,u\rangle\ .$ (25)
The second term vanishes by RobRodSad16 , Lemma 3.2. $\square$
We need the NSRI modification of this result.
  • Lemma III.2 Let $u$ be an NSRI entropy production maximizing weak solution of the Navier-Stokes equation. Then the dissipation terms $\nu_{t,Temp}$ and $\nu_{t,NL}$ are bounded in $\mathcal{H}$ with norms independent of $t$.
Proof. The proof follows that of Lemma III.1 other than the assertion that the second term on the RHS of (25) is zero. The proof ends with
$\langle\partial_{t}u,u\rangle=\nu\langle\Delta u,u\rangle-\langle(u\cdot\nabla)u,u\rangle\ .$ (26)
The dissipation terms $\nu_{t,Temp}$ and $\nu_{t,NL}$ are non-negative.
From (24), their $\mathcal{H}$ norm is constant as a function of $t$. $\square$
IV Part II: Initial Data
The continuous assumption of initial data in $\mathcal{H}$ is the main issue for Part II.
To start, we eliminate a one dimensioal space of simgular solutions.
No Global Drift
  • Proposition IV.1 Let $u$ be a general solution of the Navier-Stokes equation with smooth initial data and a stirring function $f$. Then there is a solution $\tilde{u}$ and a stirring function $\tilde{f}$ both with zero global average. $\tilde{u}$ and $\tilde{f}$ are constructed from $u$ and $f$ by a global Galilean uniform drift transformation. Thus we may always assume that $u$ and $f$ have a zero spatial average.
Proof. This is a simple computation. See also FoiManRos04 , Pg. 31 and RobRodSad16 exercise 3.8. $\square$
In the remainder of the paper, the Navier-Stokes solution is assumed to have a zero spatial average.
The global average is by itself a solution of the Navier-Stokes equation. It is the spherial harmonics $\ell=0$ mode of a general solution.
The $\ell=0$ global drift is SRI.
  • Proposition IV.1
    $u\in L^{r}(0,T;L^{s}(\mathbb{T}^{3}))$ (27)
    where $\frac{2}{r}+\frac{3}{s}=\frac{3}{2}$ for $2\leq s\leq 6$.
Proof. Recall the regularity conditions (12) and item 3 of the definition of a weak solution.
The proof is identical to the proof of RobRodSad16 , Lemma 3.5. $\square$.
  • Proposition IV.2 (i) For $u$ a weak solution of the Navier-Stokes equation,
    $\begin{split}&\int_{t_{1}}^{t_{2}}\langle\nabla u,\nabla\phi\rangle+\int_{t_{1}}^{t_{2}}\langle(u\nabla)u,\phi\rangle\\ &=\langle u(t_{1}),\phi\rangle-\langle u(t_{2}),\phi\rangle\\ \end{split}$ (28)
    for all cutoff functions $\phi$ and a.e. $t_{2}\geq t_{1}$ and a.e. $t_{1}\geq 0$ including $t_{1}=0$.
    (ii) $\partial_{t}u\in L^{4/3}(0,T;\mathcal{V}^{\prime})$
    (iii) One can modify $u$ on a set of measure zero so that $u\in\mathcal{C}([0,t];\mathcal{V}^{\prime})$ and $u$ is still a solution of its transport equation.
Proof. The proof is based on RobRodSad16 , Lemmas 3.6, 3.7. The assertion (i) follows from the choice of a cutoff function $\phi$ which is constant in $t$ on the interval $[t_{1},t_{2}]$.
To prove the assertion (ii), define $g$ by the equation
$\langle g,\phi\rangle_{\mathcal{V}^{\prime}\times\mathcal{V}}=-\langle\nabla u,\nabla\phi\rangle-\langle(u\nabla)u,\phi\rangle\ .$ (29)
Then $g$ is a bounded functional on $\mathcal{V}$ for a.e. s, because
$\begin{split}&|\langle\nabla u,\nabla\phi\rangle|\leq\|\nabla u\|\|\nabla\phi\|\;\;\mathrm{and}\\ &|\langle(u\nabla u,\phi)\rangle|\leq\|(u\nabla)u\|_{L^{6/5}}\|\phi\|_{L^{6}}\\ &\leq c\|(u\nabla)u\|_{L^{6/5}}\|\nabla\phi\|\ ,\\ \end{split}$ (30)
using the Sobolev embedding $\mathcal{H}_{1}\subset L^{6}$.
It follows that $g\in L^{4/3}(0,T;\mathcal{V}^{\prime})$. Thus also $u\in L^{4/3}(0,T;\mathcal{V}^{\prime})$.
To prove assertion (iii) and also (i), we continue to follow RobRodSad16 . By choices of test functions that are constant over time intervals, we modify $u$ on a set of measure zero so that
$u\in\mathcal{C}([0,T];\mathcal{V}^{\prime})\ .$ (31)
To prove that the modified $u$ satisfies (28) for all $t_{1},t_{2}\in(0,T]$, observe that we can choose sequences $s_{n}\rightarrow t_{1}$ and $z_{n}\rightarrow t_{t}$ such that
$\begin{split}&\int_{s_{n}}^{z_{n}}\langle\nabla u,\nabla\phi\rangle+\int_{s_{n}}^{z_{n}}\langle(u\nabla)u,\phi\rangle\\ &=\langle u(s_{n}),\phi\rangle-\langle u(z_{n}),\phi\rangle\ .\\ \end{split}$ (32)
The LHS is continuous with respect to the $s_{n}$ and $z_{n}$ limits while the RHS is continuous since $u\in\mathcal{C}([0,T];\mathcal{V}^{\prime})$. In limit as $n\rightarrow\infty$, the $t_{1},t_{2}$ transport equation (28) is satisfied for all $t_{1},t_{2}>0$. The proof with the lower limit $t_{1}=0$ is similar. $\square$
The viscous dissipation term in the above propositions is necessarily SRI due to restrictions on the possible forms of viscous dissipation. Thus there is no subscript $i$ in the viscous dissipation term.
  • Theorem IV.1 Given a velocity $u$, there is a LH solution $u(x,t)$ of the Navier-Stokes equation with this initial condition.
    The initial data is the limit as $t\rightarrow 0$ in $\mathcal{H}$ of the small $t$ values $u(x,t)$.
    Such a solution is a strong solution except for possibly isolated singular times $t_{j}\neq 0$.
Proof. The first statement is part of the LH existence proof.
According to (12) or the more detailed analysis (31), the initial conditions are achieved continuously in $\mathcal{V}^{\prime}$.
The third statement is contained in the theory of FoiManRos04 . $\square$
V Part III: Analyticity
Analyticity of solutions is the central issue of Part III.
The complexified Navier-Stokes equation is
$\frac{du}{d\zeta}+(-\nu\Delta)u+B(u)_{\mathbb{C}}=0$ (33)
FoiManRos04 , eq. (8.1). Here we write $(-\Delta)(u_{1}+iu_{2})=(-\Delta)u_{1}+i(-\Delta)u_{2}$. Due to the periodic boundary conditions, the Stokes operator has been reduced to $-\nu\Delta$. This simple identity reduces estimates of the complex term to estimates of the corresponding real terms, as in FoiManRos04 . We do not require analyticity for the Stokes operator, established in FoiManRos04 , in any case. We use the same splitting of the other complex terms in (33) into real and imaginary terms with individual estimates for the real and imaginary terms.
The complexified bilinear form $B$ is written as
$B(u,v)_{\mathbb{C}}=B(u_{1},v_{1})+B(u_{2},v_{2})+i[B(u_{1},v_{2})+B(v_{2},u_{1})]\ .$ (34)
The proof of analyticity for NSRI moments of order one is based on the energy conservation of Lemma III.2, with an entropy principle hypothesis and non-negativity of the turbulent dissipation rates. The result is a UV finite norm in $\mathcal{H}$ for $\nu_{t,NL}$ and $\nu_{t,Temp}$, with $\mathcal{H}$ continuity uniform in $t$ in a limit as the cutoffs are removed.
  • Proposition V.1 The real domain bounds of Lemma III.2 extend to the complex domain.
Proof. The real domain analysis is extended to the complex domain. As in FoiManRos04 , the complex domain analysis is repeated for the real and imaginary parts and the two bounds added. The result is still finite in $\mathcal{H}$, uniformly in $t$, with a larger but still finite norm.
For the turbulent dissipation terms, the same reduction allows a restriction to the real case only.
The proof is completed by the Lemma III.2 bounds in $\mathcal{H}$ on the turbulent dissipation terms that are uniform in $t$ as the cutoffs are removed. $\square$
Time Differentiated Analyticity
  • Proposition V.2 The time derivative of the turbulent energy $\nu_{t}$ vacuum expectation value is analytic, with a domain of analyticity at least as large as the domain of analyticity for $\nu_{t}$.
Proof. The derivative of an analytic function is analytic. $\square$
VI Part IV: Blowup
VI.1 Blowup in Finite Time
Initial conditions which are not SRI exist within the subspace of spherical harmonics spaces $\ell\geq 2$.
Blowup is considered in the space of the energy spherical harmonics expansion, so that there is no need to develop the parallel enstrophy theory. Blowup is considered within a space of finite spherical harmonics coefficients.
Making use of the assertions, the proof of blowup is valid.
The numerical program of Hou23 documents a rapid near total blowup in enstrophy, followed by a slower blowup in the energy, to complete the blowup process.
We examine only the energy spherical harmonics, so that blowup in the energy is sufficient to show existence of non-smooth solutions. Within the energy spherical harmonics, it is sufficient to restrict the statistics to the single 3D mode $\ell=2$, as blowup in any turbulent statistic shows existence of non-smooth solutions. Analyticity in time for this sequence and its time derivative was shown in Propositions V.1 and V.2.
The weak solution is chosen having a non-zero LH weak entropy maximizing solution with non-zero second moment $NSRI$. This weak solution lies in $\mathcal{H}$ and defines a LH weak solution.
This choice of initial conditions necessarily results in blowup in a finite time.
Figure 1: Two decay laws drive the blowup, for the decay of energy and the time derivative for energy, both based on analyticity. Each introduces a forbidden region, with a bound independent of time setting a minimum value. The forbidden energy region is shown in yellow across the bottom of the figure. The forbidden energy decay region is the upper triangle, shown in red. The two excluded regions overlap at a time $T^{*}$ in an excluded orange triangle. The smooth solution cannot enter the excluded triangle. It cannot continue. At $t=T^{*}$, the solution must violate its energy bound, and by a discontinuous amount. It is not smooth, and with the finite time $T^{*}$ to blow up.
  • Theorem VI.1.1 Initial conditions exist for which every entropy maximizing solution of the Navier-Stokes equation has only a finite interval of regularity.
Proof. Assume the solution is within an interval of regularity starting from its initial conditions.
Blowup occurs within the $\ell=2$ spherical harmonic space with $\nu_{t,NL}>0$. Initial conditions are chosen with this blowup mode strictly positive. These initial conditions are smooth spatially since they lie within the span of a finite number of eigenvalues of the Stokes operator.
Let $D(s)$ be the rate of decrease in this mode. The same initial conditions govern $D(s=0)$, which is also strictly positive. The domain of analyticity for $D$ is at least as large as the domain of analyticity for $\nu_{t}$. Thus $D(t)$ is strictly positive for as long as $\nu_{t}$ is strictly positive, for fixed $\ell$, $\ell\geq 2$.
While $D(s)$ is non-zero, it is bounded uniformly in $s$ by its $s$ independent positive minimum value $D_{\mathrm{min}}>0$.
Thus while non-zero, $\nu_{t}(s)$ exceeds its $s$ independent positive minimum value and decreases as a function of $s$ at the $s$ independent rate $D_{\mathrm{min}}>0$. For a finite value of $s$, $\nu_{t}(s)$ is smaller than its allowed strictly positive minimum. Thus at this time or sooner, $\nu_{t}$ falls below its smooth solution mandated minimum value, defining an endpoint to the epoch of regularity. This time is $T^{*}<\infty$. $T^{*}<\infty$ is a time of blowup.
The discontinuity (and resulting non-smoothness of the solution) occurs in the blowup mode, which is the $\ell=2$ spherical harmonic expansion coefficient for the non-linear turbulent dissipation rate $\nu_{t}$. $\square$
This set of inequalities is illustrated in Fig. 1.
  • Theorem VI.1.2 Initial conditions exist for which every entropy maximizing solution of the Navier-Stokes equation is not smooth globally in time.
Proof. We must show that the solution is not smooth at the transition point marking the end of the interval of regularity. By definition, the end of the regularity region $t_{j}$ has the property that $u(t_{j})$ is not in $\mathcal{V}$. This is equivalent to saying that $\langle\Delta u(t),u(t)\rangle$ is discontinuous at $t=t_{j}$, which demonstrates the lack of smoothness globally in time. $\square$
The energy itself, in the form of weak limits, is conserved up to viscous decay, so that weak limits of the energy are not the source of discontinuity.
The discontinuity is in the energy dissipation, as it is converted from turbulent to viscous. A finite unit of turbulent energy dissipation has transitioned through a large range of small length scales, at or below the Kolmogorov scale, where viscosity is the primary dissipative mechanism, to become totally thermal energy dissipation.
VII Acknowledgements and Disclaimer
We thank G.-Q. Chen, H. Said, T. Wallstrom and A. Rahimyar for helpful comments.
Point72 disclaimer:
The information, views, and opinions expressed herein are solely Jarret Petrillo’s and do not necessarily represent the views of Point72 or its affiliates. Point72 and its affiliates are not responsible for, and did not verify for accuracy, any of the information contained herein.