### v6

# Non-Smooth Solutions of the Navier-Stokes Equation and their Means

Abstract
Non-smooth Leray-Hopf solutions of the Navier-Stokes equation are constructed. The construction occurs in a finite periodic cube $\mathbb{T}^{3}$. Entropy production maximizing solutions with turbulent initial data are selected. The proof of finite time blowup is based on analyticity properties of the weak solutions of the Navier-Stokes equation.
The turbulent initial data is characterized in terms of its expansion in spherical harmonics basis functions.
The mean value of a weak solution of the Navier-Stokes equation is identified as a smooth solution of the Navier-Stokes equation.
1 Introduction
1.1 Main Results
The main result of this paper is the construction of non-smooth entropy production maximizing solutions of the Navier-Stokes equation of the Leray-Hopf (LH) class of weak solutions. The initial data giving rise to the non-smooth solutions is turbulent. Turbulent initial data is identified explicitly in terms of its expansion in spherical harmonics basis elements.
Initial data which is not turbulent is characterized as mean values of general Navier-Stokes LH solutions. These solutions are entropy minimizing.
The proof is divided into three main steps.
The first step demonstrates the achievability of the initial data from limits of the small time data.
The second step establishes analyticity of the weak solutions of the Navier-Stokes equation.
The third step proves blowup in finite time for selected turbulent initial conditions.
A concluding section characterizes the mean of a weak solution of the Navier-Stokes equation as a smooth solution of the Navier-Stokes equation.
In this paper, only single incompressible fluids (not mixtures) are considered. The fluids are real (not complex) valued. They are defined in a periodic spatial domain $\mathbb{T}^{3}$.
1.2 Prior Results
A review of Navier-Stokes solutions can be found in [2]. The criteria of [1] are necessary for blowup. The numerical study [21] with additional numerical and theoretical references cited provide strong evidence for the existence of non-smooth solutions of the Navier-Stokes equation.
Weak solutions for the Euler equation have been constructed, with weakly defined limits and function spaces [10, 5, 6, 16, 7, 24] and additional references cited in these sources. See [3] and references cited there for a discussion of the Navier-Stokes and Euler equations and their relation to the Onsager conjecture.
The Foias [12, 13, 33] description of Navier-Stokes turbulence as an LH weak solution provides the constructive step needed in our analysis. Other ideas, including analyticity, from this body of work are also used.
The existence of entropy production maximizing solutions of the Navier-Stokes equation is established in [18, 9]
The theory of vector and tensor spherical harmonics [38] is used.
The Lagrangian for space time smooth fluids, derived from [23, 14, 15, 20, 36, 35], has influenced our thinking. Scaling analysis, based on the renormalization group has also influenced our thinking. Of the many papers on this subject, we mention [4, 29, 40]
Additional references are added as needed throughout the text. Additional references can be found within the sources cited above. This paper is not a scholarly review.
1.3 The Fluid Millennium Prize Specification
A technical wording of this specification is given by C. Fefferman in several options. The spatially periodic option B is:
Take $\nu>0$ and $n=3$. Let $u^{0}(x)$ be any smooth, divergence-free vector field satisfying
  • smoothness, divergence free, periodicity of $u^{0}$.
  • No restrictions are placed on $f$ other than periodicity.
Then there exist smooth functions $p(x,t)$, $u_{i}(x,t)$ on $\mathbb{R}^{3}\times[0,\infty)$ that satisfy
1. the Navier-Stokes equation; 2. divergence free; 3. initial conditions given by $u^{0}(x)$; 10. periodicity; 11. smoothness.
We choose $f=0$. For initial conditions $u^{0}$ that are turbulent but otherwise unconstrained, we find a weak solution which blows up in finite time.. The prize problem is solved in the negative.
This paper and its antecedents are offered as a solution of the Millennium Fluids problem, considered as joint work of the authors contributing: J. Glimm, G.-Q. Chen, H. Said.
We show that the mean of a turbulent solution or of any weak solution is smooth. The initial data for the mean is the mean of the weak solution initial data. The mean of the weak solution data is smooth. The mean does not have general initial data. For this reason the mean is not a positive solution of the Millennium fluids problem.
2 Preliminaries
2.1 Function Spaces
Configuration Space
[…]
The bare vacuum energy of the fluid, $\langle u^{2}\rangle$, is a central concern of this paper.
Powers of the Stokes Operator
The Stokes operator is denoted $A$. For space periodic boundary conditions,
$A=-\nu\Delta\ ,$ (1)
where $\Delta$ is the Laplace operator, [13], Ch. II.6, Eq. (6.16).
The fractional power of the Stokes operator is $(-\nu\Delta)^{s}$. Using these powers, we define fractional Hilbert spaces $\mathcal{H}_{s}$ with the inner product $(u,v)_{\mathcal{H}_{s}}=(u,(-\nu\Delta)^{s}v)$. Incompressibility is imposed as part of the definition of $\mathcal{H}_{s}$.
The Spaces $\mathcal{V}$ and $\mathcal{V}^{\prime}$
The space $\mathcal{H}_{1}$ is denoted $\mathcal{V}.$ Its dual is $\mathcal{V}^{\prime}=\mathcal{H}_{-1}$.
2.2 Euler and Navier-Stokes Equations
The Euler and Navier-Stokes equations are defined physically as the law of conservation of momentum. The weak solutions of these equations are interpreted through space time integration, meaning that they are valid as an expectation value with a cutoff function $\phi$, as in (11).
The existence of nonunique and nonphysical solutions of these equations have been known for the 1D Euler equation since the days of Riemann, for the 2D Euler equation [34] and in 3D even for the Navier-Stokes equation; see [7, 8] and references cited there.
Evaluation of a divergent integral in a Sobolev space of negative index does not change the theory, but modifies the function spaces in which it is described. In view of the infinities of turbulence, these negative index Sobolev spaces are essential and can never be removed. In contrast, a cutoff, if used, is temporary and removed in the sense of limits.
The Leray Projector
Incompressibility, $\sum_{i=1}^{3}\partial_{i}u_{i}=0$, is imposed on all vector fields.
Define the Leray projector $\mathbb{P}$ as the projection onto the divergence free subspace of velocity fields, acting on an $L_{2}$ space.
$\mathbb{P}_{q}$ denotes the Leray projector acting on an $L_{q}$ space, with an adjoint $\mathbb{P}_{p}$ acting on the dual $L_{p}$ space. See [11].
The convective derivative is defined as
$\frac{D}{Dt}=\frac{\partial}{\partial t}+u\cdot\nabla\ .$ (2)
The Euler equation, expressed using Lagrangian derivatives, is
$\mathbb{P}\frac{Du_{i}}{Dt}=f_{i}\ ,$ (3)
with $f$ the stirring force, assumed divergence-free.
The Stirring Force f
Restrictions on $f$ are based on physics, as unrestricted stirring is consistent mathematically with the Euler and Navier-Stokes equations. The choice $f=0$ is made within this paper.
Stress Rates
$S_{ij}$ is the rate of stress tensor, defined as
$S_{ij}=\frac{\partial u_{i}}{\partial x_{j}}\ .$ (4)
Vorticity $\omega$ is defined as
$\omega_{i}=\nabla\times u$ (5)
and the vorticity related rate of stress is
×$\nabla\times S_{ij}=\frac{\partial\omega_{i}}{\partial x_{j}}\ .$ (6)
The stress-strain tensor
$\sigma_{ij}=\frac{1}{2}\Big{(}S_{ij}+S_{ji}\Big{)}$ (7)
specifies the fluids modeled, an incompressible isotropic Newtonian fluid.
The Navier-Stokes equation written using Lagrangian derivatives is
$\mathbb{P}\frac{Du_{i}}{Dt}=\nu\mathbb{P}(\partial_{j}\cdot S_{ij})+f_{i}\ .$ (8)
The Eulerian Navier-Stokes equation,
$\bigg{(}\frac{\partial u_{i}}{\partial t}+\mathbb{P}\Big{(}u_{j}\frac{\partial u_{i}}{\partial x_{j}}\Big{)}\bigg{)}=\nu\mathbb{P}(\nabla_{j}\cdot\sigma_{ij}(x,t))+f_{i}\ ,$ (9)
is formally equivalent to (8) with $\nu=0$ to define the Euler equation.
2.3 Cutoffs
Cutoffs are defined as a product of a spatial and a temporal cutoff. The spatial cutoff is the projection of the velocity field $u$ onto a finite dimensional space of smooth basis elements of $\mathcal{H}$. The temporal cutoff is defined by convolution by a smooth test function, which is time retarded in the sense of being supported in $t<0$.
A cutoff is defined as a modification to the theory. The cutoff theory restricts the spatial integration to a finite dimensional subspace of the configuration space $\mathcal{H}$.
The complete criteria to define a weak solution of the Navier-Stokes equation are formulated in terms of cutoffs. Cutoffs are defined as functions $\phi$ belonging to
$\begin{split}&\widetilde{\mathcal{D}}_{\sigma}(\mathbb{T}^{3})=\Big{\{}\phi\;:\;\phi=\sum_{k=1}^{N}\alpha_{k}(t)a_{k}(x),\\ &\alpha_{k}\in C_{c}^{1}([0,\infty)),\;a_{k}\in\mathcal{N},\;N\in\mathbb{N}\Big{\}}\ .\\ \end{split}$ (10)
Here $\mathcal{C}^{1}_{c}$ denotes the space of continuously differentiable functions with compact support and $\mathcal{N}$ is the set of eigenfunctions of the Stokes operator, as defined in [33], Theorem 2.24. (Caution: $\alpha_{k}\neq a_{k}$ in (10).)
Following [33], Def. 3.3, a weak solution of the Navier-Stokes equation is defined as
  1. 1.
    $u_{0}\in\mathcal{H}$
  2. 2.
    $u\in L_{\infty}(0,T;\mathcal{H})\cap L_{2}(0,T;\mathcal{H}_{1})$
  3. 3.
    $u$ satisfies weak conservation of energy with $\phi\in\widetilde{{D_{\sigma}}}(\mathbb{T}^{3})$ as test functions, according to [33], Eq. (3.1).
    $\begin{split}&\int_{0}^{s}-(u,\partial_{t}\phi)+\int_{0}^{s}\nu(\nabla u,\nabla\phi)+\int_{0}^{s}((u\cdot\nabla u),\phi)\\ &=(u_{0},\phi(0))-(u(s),\phi(s))\end{split}$ (11)
    for a.e. $s>0$.
Consequences of the definition, items 1–3, of a weak solution are given in [33].
By [33] Exercise 3.4, $u$ solves the Navier-Stokes equation weakly, relative to test functions in $\mathcal{H}_{1}$. Thus $u$ solves the Navier-Stokes equation in $\mathcal{V}^{\prime}$.
By [33], Lemma 3.4 and 3.7,
$\begin{split}&\partial_{t}u\in L^{4/3}(0,T);\mathcal{V}^{\prime})\\ &(u\cdot\nabla u)\in L^{4/3}(0,T;L^{6/5})\ .\\ \end{split}$ (12)
2.4 Second Order Phase Transitions
Decay processes, referring to the entire turbulent process and not to individual states within it, are subsumed in the generic notion of a second order phase transition, whose ultimate rest state is the hydrostatic state of a fluid at rest.
The second order phase transition occurs in four main steps, as indicated here in a formal analysis and with the first three steps studied numerically in [22], based on the analysis of DNS simulation data.
The first step is the decay of enstrophy, from a highly stirred (high vorticity) state to a state of near zero enstrophy. The second step is decay in the energy mode, driven by the cascade of energy states to smaller wavelength, where they ultimately vanish, and ending in the vacuum state of fully developed turbulence. The third step is driven by viscous dissipation, proceeding in a fourth step by radiation of thermal energy to a regime in thermal equilibrium with an ambient temperature and only thermally residual molecular motion.
The second order phase transitions for the Euler equation have the first two and fourth steps, while lacking the viscosity related third step.
2.5 Harmonic Analysis
The rotational and translation groups combine to form a semidirect product. The rotational and translation degrees of freedom of a random field, all but its absolute magnitude, are defined by its invariance properties under this semidirect product.
Fourier Analysis
Harmonic analysis of the translation invariance subgroup of the semidirect product is defined by Fourier analysis, which expands the random field into one dimensional Fourier components.
Any $u\in\mathcal{H}$, the $L_{2}$ space of turbulent fields $u$, has the Fourier expansion
$u=\sum_{k\in\mathbb{Z}^{3}}\hat{u}_{k}e^{2}\pi ix/L\ ,$ (13)
with $L$ the length of one side of the periodic domain, subject to the incompressibility condition
$k\cdot\hat{u}_{k}=0$ (14)
for all $k\in\mathbb{Z}^{3}$. Then
$L^{3}\sum_{{k\in\mathbb{Z}^{3}}}|\hat{u}_{k}|^{2}=|u|^{2}<\infty\ .$ (15)
Vector Spherical Harmonics
Spherical harmonics are described by a sum of representations of the rotation group, $O(3)$. For a scalar random field, the expansion is defined by an expansion in the basis elements $Y_{\ell m}$, with $\ell\in\mathbb{N}$ and $-\ell\leq m\leq\ell$.
Vector spherical harmonics describe the rotational degrees of freedom of a random vector field, such as a fluid velocity field. A possible spin variable to identify the particle transformation law is set to 1 for fluids. In this case, the vector spherical harmonics are described by three representations of $O(3)$. The full theory of tensor spherical harmonics is given in [38].
2.5.1 Solid Spherical Harmonics
The customary spherical harmonics, with the basis $Y_{\ell,m}$, are augmented with a magnitude $r$. The complete expansion has the basis labeled $r,\ell,m$. On restriction to the unit sphere, $r$ takes on two values, $r=\pm 1$. The plus sign selects the solid spherical harmonics solution that is regular at the origin. The minus sign selects the solid spherical harmonics solution that is regular at infinity. Detailed accounts of spherical and solid mechanics for vector and tensor fields is given in [38].
Solid mechanics harmonic analysis refers to functions that depend on $r$ as well as on the asmuthal and polar angles $\phi$ and $\theta$ which parameterize the unit sphere in $R^{3}$. [30, 32, 39, 37]. The solid mechanics solution $\Psi$, defined in terms of the parameters $(r,\theta,\phi)$ has the radial and angular separation:
$\Psi(r,\theta,\phi)=R(r)Y_{\ell,m}(\theta,\phi)\ .$ (16)
The radial part $R(r)$ has two solution classes: the spherical Bessel and Neumann functions:
$R(r)=j_{l}(kr)+y_{l}(kr)\ .$ (17)
The Neumann functions are excluded by a non-singularity condition at the origin. A singularity occurs only in the Neumann term. The general solutions are:
$\Psi_{klm}^{smooth}(r,\theta,\phi)$ $=$ $\sqrt{\frac{2}{\pi}}kj_{l}(kr)Y_{l,m}(\theta,\phi)$ (18)
$\hat{\Psi}_{klm}(r,\theta,\phi)$ $=$ $\sqrt{\frac{2}{\pi}}ky_{l}(kr)Y_{l,m}(\theta,\phi)$ (19)
Any function $f(r,\theta,\phi)$ has the Fourier expansion [39]:
[…]
The expansion of the total energy dissipation rate $\epsilon$ has a special form. The derivation [19] shows that the index $m$ does not occur in the integral,
[…]
The coefficients $C,F,S$ define the $\epsilon$ analysis. Their asymptotics relate to the regularity of $\epsilon$ in the standard way as in Fourier analysis.
Solid mechanics analysis allows a detailed evaluation of the volume integrals that determine the energy levels within a given $l$ index subspace. The combined $r,\ell,m$ analysis defines a complete expansion based on the $O(3)$ orthogonal group.
The $r$ solid harmonics expansion of a polynomial function of $u$ is independent of the spatial coordinate $x$. It is a complete description of the $t$ dependence. The $\ell,m$ expansion is independent of the time coordinate $t$ and it is a complete description of the $x$ dependence. Thus $r$ defines dynamics and $\ell,m$ define kinematics.
Both the $r$ and the $\ell,m$ variables have the dimension of energy. Their entropies are related via a Legendre transform (proven for cutoff theories) to the energy dissipation in the $r$ and $\ell$ variables.
Upon restriction to the unit sphere, the $r$ expansion takes on the values ±$\pm 1$. With this restriction $r$ is dimensionless.
Spherical harmonics extend to the vector fields labeled by energy and $\pm$ enstrophy. This extension applies to solid harmonics, with additional labels for $\Psi$ and $\hat{\Psi}$ to describe the solid mechanics solution.
Fixed Time and Fixed Space Harmonics
The functions $Y_{\ell,m}$ provide a complete orthonormal basis for the spatial (fixed time) $L_{2}$ space, with the restriction $r=1$. The result is a function of the three spatial variables.
General functions are then a product of the fixed space functions and a remainder.
The remainder is the $r$ expansion, defined at a fixed spatial location. It is a function of time. The remainder is described by $r$ as a configuration space variable and by a conjugate momentum $\pi$. $\pi$, defined at a fixed space location, is a function of $x_{1},x_{2},x_{3}$ equivalence classes. $\pi$ thus has a single (generic) spatial variable, which does not refer to any individual $x_{1},x_{2},x_{3}$.
2.5.2 Quotient Spaces
We define a quotient space by identifying all fixed time harmonic expansion terms. The quotient space is isomorphic to the fixed space harmonic expansion.
The fixed time spherical harmonics are defined by restriction of all variables to be spatial only. The result is a function of three spatial variables. From the point of view of the fixed time expansion, $r$ is a scalar. $r$ is eliminated on restriction to the sphere with $r=\pm 1$ and selection of the solution regular at the origin, with $r=1$. At fixed time, all variables are in configuration space, which is defined as functions of 3 spatial variables.
The fixed space spherical harmonics are defined with one temporal variable, a configuration space variable.
In the $r$ expansion, $r$ is the one dimensional configuration space variable. Associated with this variable is its conjugate momenta $\pi$, which is a stochastic function of a one dimensional variable.
Since $r$ and $\pi$ are defined at fixed space variables, the spatially dependent $\pi$ is defined over permutation equivalence classes of spatial variables. $\pi$ is thus a function of time and a generic (1D) spatial variable. The spatial variable has no reference to specific $x_{i},x_{2},x_{3}$ spatial coordinates. $r$ and $\pi$ define loop variables. The loop variables provide a complete basis for the $r,\pi$ expansion, defined at a fixed spatial location.
Young’s Inequality
Young’s inequality [13], eq. (4.17), pg. 20. states that
$ab\leq\frac{a^{p}}{p}+\frac{b^{p^{\prime}}}{p^{\prime}}$ (22)
for all $a,b>0$ and for $p,p^{\prime}$ Holder conjugate indices.
Young’s inequality can be used to bound cross terms, so that the details of the quotient spaces can by bypassed with separate estimates of the fixed space and fixed time cases.
2.6 Navier-Stokes Term by Term Regularity
If $u$ is a weak solution of the Navier-Stokes equation, then
  1. 1.
    $u\in L^{\infty}(0,T;\mathcal{H})\bigcap L^{2}(0,T;\mathcal{V})$,
  2. 2.
    $\partial_{t}u\in L^{4/3}(0,T;\mathcal{V}^{\prime})$,
  3. 3.
    $(u\cdot\nabla)u\in L^{4/3}(0,T;L^{6/5})\subset L^{4/3}(0,T;\mathcal{V}^{\prime})$,
  4. 4.
    $u\in C(0,\infty,\mathcal{V}^{\prime})$ with $C$ denoting the space of continuous functions
  5. 5.
    $\partial_{t}u-\Delta u+(u\cdot\nabla)u=0\in\mathcal{V}^{\prime}$ as an element of $L^{4/3}(0,T;\mathcal{V}^{\prime})$
by [33], Lemma 2 (with following discussion) for item 1, Lemma 3.7 for item 2, Lemma 3.4 for item 3 and Definition 3.1. Item 4 is a consequence of item 3 and item 5 is a consequence of items 1 – 3.
  • Proposition 2.6.1 The terms in Item 5 are the terms of the Navier-Stokes equation. Each term is individually spatially in $\mathcal{V}^{\prime}$ and temporally in $L^{4/3}$.
Proof. As stated above. $\square$
2.7 The Schwartz Space
The Schwartz space $\mathcal{S}$ of smooth vector fields is defined as
$\mathcal{S}=C^{\infty}(0,T;C^{\infty}(\mathbb{T}^{3}))$ (23)
with its customary topology, and its dual space is
$\mathcal{S}^{\prime}=C^{\infty^{\prime}}(0,T;C^{\infty^{\prime}}(\mathbb{T}^{3}))$ (24)
$\mathcal{S}$ and $\mathcal{S}^{\prime}$ are closed under multiplication, so that polynomials of elements of $\mathcal{S}$ are also in $\mathcal{S}$ and polynomial elements of $\mathcal{S}^{\prime}$ are also in $\mathcal{S}^{\prime}$.
2.8 Dissipation Operators
The various forms of turbulent and viscous dissipation are summarized by (25).
$\begin{split}&\nu_{t,NL}=\langle\sum_{j}u_{j}\partial_{j}u_{i}\rangle\\ &\nu_{t,Temp}=-\langle u_{i},\partial_{t}u_{k}\rangle\\ &\nu_{t}=\nu_{t,NL}+\nu_{t,Temp}\\ &\nu_{Tot}=\nu_{t,Temp}+\nu_{t,NL}+\nu\ .\\ \end{split}$ (25)
Within the solid mechanics expansion variables, each of the $r,\ell,m$ expansions are separately evaluated by the solid harmonics solution which is regular at the origin. In this way, each of the individual terms in (25) is nonnegative.
In estimating polynomials of the dissipation operators (25), cross terms are bounded by Young’s inequality, so that bounds of polynomials in the individual terms are sufficient.
3 The Entropy Principle
The entropy principle as used in this paper states that the solution $u(x,t)$ of the Navier-Stokes equation, maximizes the production of entropy relative to other weak solutions of the Navier-Stokes equation. This principle is applied within the space of solid spherical harmonic expansion coefficients. It is convenient to maximize individually the entropy in both the $r,\pi$ and the $\ell,m$ parameters.
The entropy maximizing state is computed as the equally weighted sum for the $r,\pi$ and $\ell,m$ values.
See [18, 9] for a related discussion. Uniqueness of the entropy maximizing state will be addressed in a future publication with G.-Q. Chen and H. Said.
It is necessary to specify the physical laws assumed for the current analysis. Assuming isothermal, single fluid physics, entropy is time independent, while the energy satisfies a heat equation with decay of energy.
Other physical assumptions, such as thermal physics and compressibility allow time dependent entropy.
For the isothermal physics considered here, the entropy has a dependence on the constant temperature of the solution. A convenient normalization is to set the constant temperature to unity.
Positivity of the $\nu_{t}$ Variables
  • Proposition 3.1 Entropy maximizing dissipation operators $\nu_{t,Temp}$ and $\nu_{t,NL}$ are non-negative in their energy dissipation.
Proof. We treat entropy maximizing and energy dissipation maximizing as identical. The maximum rate of energy dissipation chooses a positive sign for the dissipation. $\square$
A weak form of the (exact) conservation of energy is contained in [26], with an elegant and simple proof. We note the relation of this proof to earlier studies of [25, 28, 17]. The proof assumes a sum over repeated indices (SRI). We adapt it to the case NSRI (which is defined as no SRI). It will play a key role in the paper.
The result of [27] is formulated as a Lemma 1:
  • Lemma 3.1 Let $u$ be a weak solution of the Navier-Stokes equation. Then
    $\langle\partial_{t}u,u\rangle=-\|\nabla u\|^{2}$ (26)
    for a.e. $t$, where the brackets denote a dual pairing between $\mathcal{V}^{\prime}$ and $\mathcal{V}$.
Proof. We have already observed that the Navier-Stokes equation is satisfied for a.e. $t$ while $u\in\mathcal{V}$. We take the inner product with $u$ and vacuum expectation values to obtain
$\langle\partial_{t}u,u\rangle=\nu\langle\Delta u,u\rangle-\langle(u\cdot\nabla)u,u\rangle\ .$ (27)
The second term vanishes by [33], Lemma 3.2. $\square$
We need the NSRI modification of this result.
  • Lemma 3.2 Let $u$ be an NSRI entropy production maximizing weak solution of the Navier-Stokes equation. Then the dissipation terms $\nu_{t,Temp}$ and $\nu_{t,NL}$ are bounded in $\mathcal{H}$ with norms independent of $t$.
Proof. The proof follows that of Lemma 3.1 other than the assertion that the second term on the RHS of (27) is zero. The proof ends with
$\langle\partial_{t}u,u\rangle=\nu\langle\Delta u,u\rangle-\langle(u\cdot\nabla)u,u\rangle\ .$ (28)
The dissipation terms $\nu_{t,Temp}$ and $\nu_{t,NL}$ are non-negative.
From (26), their $\mathcal{H}$ norm is constant as a function of $t$. $\square$
4 Initial Data
No Global Drift
  • Proposition 4.1 Let $u$ be a general solution of the Navier-Stokes equation with smooth initial data and a stirring function $f$. Then there is a solution $\tilde{u}$ and a stirring function $\tilde{f}$ both with zero global average. $\tilde{u}$ and $\tilde{f}$ are constructed from $u$ and $f$ by a global Galilean uniform drift transformation. Thus we may always assume that $u$ and $f$ have a zero spatial average.
Proof. This is a simple computation. See also [13], Pg. 31 and [33] exercise 3.8. $\square$
In the remainder of the paper, the Navier-Stokes solution is assumed to have a zero spatial average.
The global average is by itself a solution of the Navier-Stokes equation. It is the spherical harmonics $\ell=0$ mode of a general solution.
The $\ell=0$ global drift is SRI.
  • Proposition 4.1
    $u\in L^{r}(0,T;L^{s}(\mathbb{T}^{3}))$ (29)
    where $\frac{2}{r}+\frac{3}{s}=\frac{3}{2}$ for $2\leq s\leq 6$.
Proof. [33], Lemma 3.5. $\square$.
  • Proposition 4.2 (i) For $u$ a weak solution of the Navier-Stokes equation,
    $\begin{split}&\int_{t_{1}}^{t_{2}}\langle\nabla u,\nabla\phi\rangle+\int_{t_{1}}^{t_{2}}\langle(u\nabla)u,\phi\rangle\\ &=\langle u(t_{1}),\phi\rangle-\langle u(t_{2}),\phi\rangle\\ \end{split}$ (30)
    for all cutoff functions $\phi$ and a.e. $t_{2}\geq t_{1}$ and a.e. $t_{1}\geq 0$ including $t_{1}=0$.
    (ii) $\partial_{t}u\in L^{4/3}(0,T;\mathcal{V}^{\prime})$
    (iii) One can modify $u$ on a set of measure zero so that u is a continuous function of $t$ with values in $\mathcal{V}^{\prime})$ and $u$ is still a solution of its transport equation.
Proof. The proof is based on [33], Lemmas 3.6, 3.7. $\square$
The viscous dissipation term in the above propositions is necessarily SRI due to restrictions on the possible forms of viscous dissipation. Thus there is no subscript $i$ in the viscous dissipation term.
  • Theorem 4.1 Given a velocity $u$ with no global drift, there is a LH solution $u(x,t)$ of the Navier-Stokes equation with this initial condition.
    The initial data is the limit as $t\rightarrow 0$ in $\mathcal{H}$ of the small $t$ values $u(x,t)$.
    Such a solution is a strong solution except for possibly isolated singular times $t_{j}\neq 0$.
Proof. The first statement is part of the LH existence proof.
According to (12), the initial conditions are achieved continuously in $\mathcal{V}^{\prime}$.
As is conventional in the study of the Navier-Stokes equation, a global spatial average is removed by a Galilean transformation. With the subtraction of this global spatial average, the initial conditions are achieved in $\mathcal{H}$.
The third statement is contained in the theory of [13]. $\square$
5 Analyticity
The complexified Navier-Stokes equation is
$\frac{du}{d\zeta}+(-\nu\Delta)u+B(u)_{\mathbb{C}}=0$ (31)
[13], eq. (8.1). Here we write $(-\Delta)(u_{1}+iu_{2})=(-\Delta)u_{1}+i(-\Delta)u_{2}$. Due to the periodic boundary conditions, the Stokes operator has been reduced to $-\nu\Delta$. This simple identity reduces estimates of the complex term to estimates of the corresponding real terms, as in [13]. We do not require analyticity for the Stokes operator, established in [13], in any case. We use the same splitting of the other complex terms in (31) into real and imaginary terms with individual estimates for the real and imaginary terms.
The complexified bilinear form $B_{\mathbb{C}}$ is written as sum of its real and imiganary parts
The proof of analyticity for NSRI moments of order one is based on the energy conservation of Lemma 3.2, with an entropy principle hypothesis and non-negativity of the turbulent dissipation rates. The result is a UV finite norm in $\mathcal{H}$ for $\nu_{t,NL}$ and $\nu_{t,Temp}$, with $\mathcal{H}$ continuity uniform in $t$ in a limit as the cutoffs are removed.
  • Proposition 5.1 The real domain bounds of Lemma 3.2 extend to the complex domain.
Proof. The real domain analysis is extended to the complex domain. As in [13], the complex domain analysis is repeated for the real and imaginary parts and the two bounds added. The result is still finite in $\mathcal{H}$, uniformly in $t$, with a larger but still finite norm.
For the turbulent dissipation terms, the same reduction allows a restriction to the real case only.
The proof is completed by the Lemma 3.2 bounds in $\mathcal{H}$ on the turbulent dissipation terms that are uniform in $t$ as the cutoffs are removed. $\square$
Time Differentiated Analyticity
  • Proposition 5.2 The time derivative of the turbulent energy $\nu_{t}$ vacuum expectation value is analytic, with a domain of analyticity at least as large as the domain of analyticity for $\nu_{t}$.
Proof. The derivative of an analytic function is analytic. $\square$
6 Blowup
  • Definition 6.1 Initial data $u_{0}$ is called non-turbulent (or smooth) if its expansion in spherical harmonics basis functions is proportional to a mean value in the $m$ variables and a mean value in the $\ell$ variables.
    The data is otherwise turbulent. The data is strictly turbulent if it is orthogonal to the smooth data at each fixed $m$ and also orthogonal to the smooth data as a function of $\ell$.
The three dimensional spherical harmonics space defined by $\ell=1$ is used to illustrate this definition. In this space, the mean is the lowest energy state in this subspace. It is SRI. The two states orthogonal to it are strictly turbulent. This two dimensional space will be used to construct blowup solutions. This space is NSRI.
Any finite dimensional subspace is not orthogonal to the $\ell$ mean and is thus turbulent.
Making use of the assertions, the proof of blowup is will be shown to be valid.
The numerical program of [21] documents a rapid near total blowup in enstrophy, followed by a slower blowup in the energy, to complete the blowup process.
The weak solution is chosen as LH weak entropy maximizing solution with strictly turbulent initial conditions. This weak solution lies in $\mathcal{H}$ and defines a LH weak solution.
This choice of initial conditions is conjectured to result in blowup in a finite time. Restricted to $\ell=1$, we prove blowup.
Figure 1: Two decay laws drive the blowup, for the decay of energy and the time derivative for energy, both based on analyticity. Each introduces a forbidden region, with a bound independent of time setting a minimum value. The forbidden energy region is shown in yellow across the bottom of the figure. The forbidden energy decay region is the upper triangle, shown in red. The two excluded regions overlap at a time $T^{*}$ in an excluded orange triangle. The smooth solution cannot enter the excluded triangle. It cannot continue. At $t=T^{*}$, the solution must violate its energy bound, and by a discontinuous amount. It is not smooth, and with the finite time $T^{*}$ to blow up.
  • Conjecture 6.1 Each entropy maximizing solution of the Navier-Stokes equation with strictly turbulent initial conditions has only a finite interval of regularity. It is not globally smooth in time.
  • Theorem 6.1 Each entropy maximizing solution of the Navier-Stokes equation with strictly turbulent initial conditions in the spherical harmonics subspace $\ell=1$ has only a finite interval of regularity. It is not globally smooth in time.
Proof. Assume the solution is within an interval of regularity starting from its initial conditions.
Blowup occurs within the $\ell=1$ spherical harmonic space with $\nu_{t,NL}>0$. Initial conditions are chosen with this blowup mode strictly positive. These initial conditions are smooth since they lie within the span of a finite number of eigenvalues of the Stokes operator.
Let $D(s)$ be the rate of decrease in this mode. The same initial conditions govern $D(s=0)$, which is also strictly positive. The domain of analyticity for $D$ is at least as large as the domain of analyticity for $\nu_{t}$. Thus $D(t)$ is strictly positive for as long as $\nu_{t}$ is strictly positive, for fixed $\ell$, $\ell\geq 2$.
While $D(s)$ is non-zero, it is bounded uniformly in $s$ by its $s$ independent positive minimum value $D_{\mathrm{min}}>0$.
Thus while non-zero, $\nu_{t}(s)$ exceeds its $s$ independent positive minimum value and decreases as a function of $s$ at the $s$ independent rate $D_{\mathrm{min}}>0$. For a finite value of $s$, $\nu_{t}(s)$ is smaller than its allowed strictly positive minimum. Thus at this time or sooner, $\nu_{t}$ falls below its smooth solution mandated minimum value, defining an endpoint to the epoch of regularity. This time is $T^{*}<\infty$. $T^{*}<\infty$ is a time of blowup.
The discontinuity (and resulting non-smoothness of the solution) occurs in the blowup mode, which is the $\ell=1$ spherical harmonic expansion coefficient for the non-linear turbulent dissipation rate $\nu_{t}$.
We next show that the solution is not smooth at the transition point marking the end of the interval of regularity. By definition, the end of the regularity region $t_{j}$ has the property that $u(t_{j})$ is not in $\mathcal{V}$. This is equivalent to saying that $\langle\Delta u(t),u(t)\rangle$ is discontinuous at $t=t_{j}$, which demonstrates the lack of smoothness globally in time. $\square$
This set of inequalities is illustrated in Fig. 1.
The energy itself, in the form of weak limits, is conserved up to viscous decay, so that weak limits of the energy are not the source of discontinuity.
The discontinuity is in the energy dissipation, as it is converted from turbulent to viscous. A finite unit of turbulent energy dissipation has transitioned through a large range of small length scales, at or below the Kolmogorov scale, where viscosity is the primary dissipative mechanism, to become totally viscous energy dissipation.
7 The Mean
The mean $\mathcal{M}(u)$, defined in the vector spherical harmonic space $\mathcal{H}$ is a sum of three separate scalar spherical harmonics means. One is defined in the energy expansion and the other two are the $\pm$ oriented enstrophy expansions. The $\pm$ vorticity orientation reduces to two separate formulas for the mean $\mathcal{M}(u)$. The energy expansion has no dependance on vorticity, so that the energy mean can be determined independently of the spherical harmonics $m$ values.
7.1 Entropy Production Minimization
From the solid harmonics expansion of a weak solution $u$ of the Navier-Stokes equation or of any polynomial function of $u$, we note that the total entropy is equal to the sum of two individual entropies, defined by the $r$ and the $\ell$ solid harmonics expansion variables.
A Legendre transform relates $\ell$ energy dissipation to the $\ell$ entropy.
All proofs are of the form of uniform properties of cutoff quantities. The Legendre transforms are established in the cutoff case. Thus we can reason in terms of energy dissipation.
The Minimum Entropy Principle
The minimum entropy principle (when valid for a specific solution) states that a nonnegative polynomial function of the solution has a minimum rate of entropy production relative to the same polynomial function of any other weak solution of the Navier-Stokes equation with the same initial conditions.
A Key Distinction
The key distinction between the two (opposite) cases of entropy production maximizing vs. minimizing is the fact that for an entropy production minimizing solution, the turbulent dissipation rate $\nu_{t}$ is zero.
The blowup in Fig. 1 changes its character: the diagonal sloping line generated by the time derivative is now horizontal, and never intersects the horizontal line below it defined by the energy. Thus there is no end to the duration of the strong solution, and the mean solution is smooth for all space and time.
  • Remark 7.1.1 The minimum entropy principle is used in this paper to study Navier-Stokes solutions and also nonnegative polynomial functions of $\nu_{t,NL}$ and $\nu_{t,Temp}$.
    The spherical harmonics energy defined by $\ell$ is intrinsically nonnegative. A nonnegative entropy principle polynomial for $\ell$ must have nonnegative even coefficients and zero odd coefficients.
  • Theorem 7.1.1 Any element of $\mathcal{H}$ is expressed uniquely in terms of its spherical and solid harmonics expansion variables, $r,\ell,m$.
    The mean $\mathcal{M}(u)$ is the sum (normalized by division by the number of terms in the sum) of
    1. 1.
      the $r$ configuration space (magnitude equal ±$\pm 1$) mean,
    2. 2.
      the loop expansion phase space mean defined at fixed spatial coordinates,
    3. 3.
      the $\ell$ terms defined at fixed $t$,
    4. 4.
      the $m$ mean defined at fixed $t$ and fixed $\ell$.
    The mean has entropy zero.
The proof is placed in Sec. 9.
7.2 Regularity of the Mean
For a nonnegative polynomial function $p(u)$ of a weak solution of the Navier-Stokes equation, its average value $\mathcal{M}(|p(u)|)$ is defined as the ratio $\langle|p(u)|\rangle/\langle 1\rangle$. In simple terms it is the sum of values divided by the number of terms in the sum.
Here we study $\mathcal{M}(\nu_{t,Temp})$ and $\mathcal{M}(\nu_{t,NL})$. Products of distributions are not a concern, as we introduce cutoffs (so that all distributions become smooth functions which can be multiplied) and then prove results that are uniform as the cutoffs are removed.
The complete mean is the sum of the means defined by the spherical harmonic $r,\ell$ means.
  • Theorem 7.2.1
    $\mathcal{M}(\nu_{t,Temp})=0\in L^{4/3}(0,T;\mathcal{V}^{\prime})\subset\mathcal{S}^{\prime}$ (32)
    $\mathcal{M}(\nu_{t,NL})=0\in L^{4/3}(0,T;\mathcal{V}^{\prime})\subset\mathcal{S}^{\prime}$ (33)
  • Theorem 7.2.2
    $\nu_{t,Temp}(\mathcal{M}(\nu_{t,Temp}))=0\in\mathcal{S}^{\prime}$ (34)
    $\nu_{t,NL}(\mathcal{M}(\nu_{t,NL}))=0\in L^{4/3}(0,T;\mathcal{H}_{-3})\subset{S}^{\prime}\ .$ (35)
Intuitively, the evaluation is zero because the turbulent fluctuation of a turbulent fluctuation is zero.
Proofs are placed in Sec. 9.
8 The Mean is a Smooth Solution of the Navier-Stokes Equation
The main result of this section is the construction of smooth solutions of the Navier-Stokes equation with smooth non-turbulent initial data. The smooth solution is the mean of a weak solution. Its entropy is zero, so that it can also be characterized as an entropy production minimizing solution. We show that
  1. 1.
    Initial conditions: The mean of a weak solution of the Navier-Stokes equation with initial data $u^{0}\in\mathcal{H}(\mathbb{T}^{3})$ (the space of $L_{2}$ velocity fields defined on the cube $\mathbb{T}^{3}$ with periodic boundary conditions) the initial conditions are assumed in a continuous fashion, in the sense that
    $\mathcal{M}(u^{0})=\lim_{t\rightarrow 0}\mathcal{M}(u(t))\ .$ (36)
    in the topology of $\mathcal{H}(\mathbb{T}^{3})$.
    All vector fields are assumed to be divergence-free.
  2. 2.
    The mean of a weak solution of the Navier-Stokes equation is a strong solution. It is smooth for all time.
  • Theorem 8.1 Assume smooth initial conditions $u^{0}\in C^{\infty}$.
    The mean is smooth.
    The mean assumes its initial conditions (36).
    The mean of a weak solution $u$ of the Navier-Stokes equation in $\mathcal{H}$ is a solution of the Navier-Stokes equation in $\mathcal{H}$.
Proof. By Theorem 7.2.1, $\mathcal{M}(\nu_{t,NL}(u))=0\in\mathcal{H}_{-2}\subset\mathcal{S}^{\prime}$. and $\mathcal{M}(\nu_{t,Temp}(u))=0\in\mathcal{H}_{-2}\subset\mathcal{S}^{\prime}$.
To find the equation which the mean satisfies, we reason directly. $u$ satisfies the equation
$\partial_{t}u+\nu_{t,Temp}+\nu_{t,NL}(u)=\nu\Delta u\in\mathcal{S}^{\prime}\ .$ (37)
Since the mean $\mathcal{M}$ is linear, it is applied to each of the terms of (37) individually, in the space $\mathcal{V}^{\prime}\subset\mathcal{S}^{\prime}$. By Theorem 7.1.2 the mean is a solution of the Navier-Stokes equation in $\mathcal{V}^{\prime}$. $\mathcal{M}(\nu_{t,NL}(u))$ and $\mathcal{M}(\nu_{t,Temp}(u))$ are each the 0 element of $\mathcal{S}^{\prime}$.
By Theorem 7.2.1, there remains the equation
$\partial_{t}\mathcal{M}(u)=\nu\Delta\mathcal{M}(u)\ ,$ (38)
is valid in $\mathcal{V}^{\prime}\subset\mathcal{S}^{\prime}$.
This is the equation is also the heat equation. It is well known that the solutions in $\mathcal{S}^{\prime}$ of the heat equation with $\mathcal{S}$ initial conditions are smooth, belonging to $\mathcal{S}$, to prove the first assertion.
Solutions of the heat equation with initial data $u^{0}\in\mathcal{H}$ satisfy continuity of initial data (36). This statement is applied not to the weak solution $u$ but to $\mathcal{M}(u)$, with assumed unrestricted smooth initial data $\mathcal{M}(u^{0})$. Thus we assume $\mathcal{M}(u^{0})\in\mathcal{S}$, and the initial conditions (36) are assumed, the second assertion.
To prove the third assertion, recall the 0 evaluation in $\mathcal{S}^{\prime}$ of $\nu_{t,NL}(\nu_{t,NL}(\mathcal{M}{)u})$ and of $\nu_{t,Temp}(\mathcal{M}(\nu_{t,Temp}))$ according to Theorem 7.2.2.
Thus (38) is the Navier-Stokes equation in $\mathcal{S}^{\prime}$.
As inferred above, $\mathcal{M}(u)\in\mathcal{S}$ is smooth, so that (38) is the Navier-Stokes equation in $\mathcal{S}$. $\square$.
9 Appendix
9.1 Theorem 7.1.1
  • Theorem 9.1.1 The minimum entropy solution of the Navier-Stokes equation is unique. It is an element of $\mathcal{V}^{\prime}$.
Proof: By items 2,3, $\nu_{t,Temp}\in\mathcal{V}^{\prime}$ and $\nu_{t,NL}\in\mathcal{V}^{\prime}$.
By general theories of positive definite forms, uniqueness of a minimum is a result of strict convexity.
Strict convexity of the energy dissipation is a result of the viscous dissipation term $\nu\Delta u$, which sets a strictly positive lower bound on the dissipation as $\nu\Delta u$, with a gradient in $\mathcal{V}^{\prime}$. This completes the proof for a solution in $\mathcal{V}^{\prime}$. $\square$
Proof of Theorem 7.1.1 A mean is not a random variable. It is the expectation value of a random variable.
The first statement is elementary.
The mean of any of the summands 1 – 4 in the statement of the theorem is defined as
$\frac{\sum\mathrm{summand}}{\mathrm{number}\;\mathrm{of}\;\mathrm{terms}}\ .$ (39)
In the cutoff cases where this formula is used, the sum contains only a finite number of terms.
The $r$ phase space entropy production and $r$ phase space energy dissipation are evaluated at fixed spatial coordinates while the $\ell$ entropy production and energy dissipation are evaluated with the time coordinate fixed.
The $r$ configuration space dissipation is a sum of the energy dissipation for the two solid harmonic solutions, regular and irregular at the origin. The solution irregular at the origin is chosen as minimizing the energy dissipation. Uniqueness of this (one dimensional) minimum dissipation state follows.
The $r$ conjugate momentum space is the expansion space of the conjugate momenta $\pi$, called the loop expansion. Uniqueness of the entropy minimizing loop expansion variable is a consequence of Theorem 9.1.1.
The volume of the entropy minimizing solution is its $L_{1}$ expectation value, also equal to its $L_{1}$ norm. The volume of the solution divided by its $L_{1}$ norm is 1. With the unity temperature normalization, the entropy is the $L_{1}$ norm of this ratio.
By the definition of the entropy, the $r$ entropy of the mean is $\log 1=0$.
Next we specialize to a fixed time, so that the spherical harmonics provides a complete expansion of the mean of the $\ell,m$ variables.
The denominator in (39) is the number of energy dissipation terms in the numerator. We consider the spherical harmonic basis $Y_{\ell,m}$ and a partial mean $\mathcal{M}_{\ell}$ defined in terms of the sum over $\ell$.
All steps as in the $r$ case can be followed. The energy dissipation is intrinsicly nonnegative, to reduce the solution space to one dimension. To arrive at a one dimensional space, we use Theorem 7.1.2, which states that the $\ell$ entropy minimum is unique, or equivalently, that the $\ell$ minimum energy dissipation state is unique.
(Related analysis focused on entropy production maximization will be presented in joint work with H. Said and G.-Q. Chen.)
The conclusion is then the same as in the $r$ case. The entropy of the $\ell$ mean $\mathcal{M}_{\ell}(u)$ is zero.
The plus sign is chosen since the vacuum expectation value is nonnegative. The rest of the proof follows as in the $r$ case.
Thus
$\langle M(u)\rangle=\langle\mathcal{M}_{\ell}(u)\rangle_{\mathcal{H}}=1\ .$ (40)
Since the volume of $\langle\mathcal{M}_{\ell}(u)\rangle_{\mathcal{H}}=1$, the entropy (the log of this volume) of $\mathcal{M}_{\ell}=0$.
Entropy is always 0 or larger. Thus the mean is entropy minimizing. $\square$
  • Corollary 9.1.1 The mean satisfies the minimum entropy principle.
Proof. Since $\mathcal{S}$ and $\mathcal{S}^{\prime}$ are closed under multiplication, the polynomial powers needed here are automatic. Thus the proof follows as in Theorem 7.1.1. $\square$
9.2 Proof of Theorem 7.2.1
By items 2,3, $\nu_{t,Temp}\in\mathcal{V}^{\prime}$ and $\nu_{t,NL}\in\mathcal{V}^{\prime}$. The mean of $\nu_{t,Temp}$ is at least as regular as $\nu_{t,Temp}$. Likewise, the mean of $\nu_{t,NL}$ is at least as regular as $\nu_{t,NL}$.
To justify the value 0 for the mean, consider item 2 of Sec. 2.6 for (32) and item 3 for (33).
The time derivative $\partial_{t}u$ occurs exclusively in the $r$ expansion. We assert that the $r$ terms contributing to $\mathcal{M}(\nu_{t,Temp})$ sum to zero in $\mathcal{V}^{\prime}$.
The nonlinear term $-\langle u_{i}\partial_{t}u_{k}\rangle$ occurs only in the $r$ solid harmonics expansion. The term is evaluated at fixed spatial coordinates. As a result of Lemma 3.2, the energy in the $r$ loop expansion is conserved up to the amount dissipated by the viscous term $\Delta u$. This conservation is preserved by taking a mean. Thus the $r$ loop expansion terms contributing to $\mathcal{M}(\nu_{t,temp})$ sum to zero in $\mathcal{V}^{\prime}$.
The $r$ term (32) selects the solid spherical harmonics configuration space solution which is regular at the origin.
The proof of (32), based on the $r$ configuration space and the $r$ loop expansion is complete.
Next, we prove (33).
The nonlinear term $(u\cdot\nabla)u$ occurs exclusively in the $\ell,m$ spherical harmonics expansion. This expansion is applied to the solution considered at a fixed time. At fixed time, the time derivative and the $r$ expansion are removed from item 3 of Sec. 2.6. As a result of Lemma 3.2 the energy (in the $\ell,m$) expansion is conserved exactly up to the amount dissipated by the viscous term $\Delta u$. This conservation is preserved by taking a mean.
Assuming this, the $\ell,m$ terms contributing to $\mathcal{M}(\nu_{t,NL})$ sum to zero.
This term is nonnegative, which was to be demonstrated. The quadratic polynomial $\nu_{t,NL}$ has a nonnegative quadratic coefficient, to complete the proof of (33). $\square$
9.3 Proof of Theorem 7.2.2
To prove (34) analyticly, consider
$\begin{split}&\nu_{t,Temp}(\mathcal{M}(\nu_{t,Temp}))=\nu_{t,Temp}[\mathcal{M}\langle(-u_{i}\partial_{t}u)\rangle]\\ &=\langle\langle u_{j}\partial_{t}u\rangle\mathcal{M}(u_{i}\partial_{t}u)\rangle\\ &=\langle u_{j}\partial_{t}u\rangle\langle\mathcal{M}(u_{i}\partial_{t}u)\rangle=0\ .\\ \end{split}$ (41)
Next we prove (35) analytically. Consider
$\begin{split}&\nu_{t,NL}\mathcal{M}(\nu_{t,NL})=((u\cdot\nabla)\cdot\nabla u)\mathcal{M}((u\cdot\nabla)u)\\ &((u\cdot\nabla)\cdot\nabla u)((u\cdot\nabla)(u\cdot\nabla)u)\ .\\ \end{split}$ (42)
and focus on the RHS two rightmost $\nabla$. The first and second of the four $\nabla$ describe fluctuations in the rightmost third and fourth $\nabla$.
Direct evaluation of these two rightmost $\nabla$ of (LABEL:eq:mean-mean) identifies the result as the zero element in the space of of cubic polynomials in elements of the space $L^{4/3}(0,T;\mathcal{H}_{-2})$.
The left two $\nabla$, acting on this zero element, is the zero element in the space of quintic polynomials in the still more singular space $L^{4/3}(0,T;\mathcal{H}_{-4})$. Thus $\nu_{t,NL}(\nu_{t,NL})=0\in\mathcal{S}^{\prime}$. This completes the proof of (35). $\square$
10 Acknowledgements and Disclaimer
We thank G.-Q. Chen, H. Said, T. Wallstrom and A. Rahimyar for helpful comments.
Point72 disclaimer:
The information, views, and opinions expressed herein are solely Jarret Petrillo’s and do not necessarily represent the views of Point72 or its affiliates. Point72 and its affiliates are not responsible for, and did not verify for accuracy, any of the information contained herein.